# New determination of the neutrino hadronic production cross sections from GeV to beyond PeV energies

### v1

@@ VI Discussion and conclusions
In this paper, we present a new evaluation of the neutrino production cross section from $p+p$ collisions, building on our previous analysis of the production cross section for $\pi^{\pm}$, $K^{\pm}$, and minor baryonic and mesonic channels that contribute at least at the 0.5% level. Cross sections for scatterings involving nuclei heavier than protons are also derived. The novelty of our approach, compared to the state of the art that relies on Monte Carlo predictions, is that we shape the $\sigma_{\rm inv}$ analytically, and directly fit the model to the available data. We work with neutrino energies from $10$ GeV to $10^{7}$ GeV, and correspondingly to incident proton (nuclei) energies from $10$ GeV to $10^{9}$ GeV (GeV/n). Our results include a realistic and conservative estimate of the uncertainties affecting the differential cross section $d\sigma/dE_{\nu}$, understood as the sum over all production channels. The uncertainty is estimated to be $5\%$ for $E_{\nu}\leq 1000$ GeV, increasing to $10\%$ at higher energies in the TeV range and reaching 20% at $E_{\nu}\simeq 10^{7}$ GeV.

### v2

@@ VI Discussion and conclusions
In this paper, we present a new evaluation of the neutrino production cross section from $p+p$ collisions, building on our previous analysis of the production cross section for $\pi^{\pm}$, $K^{\pm}$, and minor baryonic and mesonic channels that contribute at least at the 0.5% level. Cross sections for scatterings involving nuclei heavier than protons are also derived. The novelty of our approach, compared to the state of the art that relies on Monte Carlo predictions, is that we shape the $\sigma_{\rm inv}$ analytically, and directly fit the model to the available data. While the parameterization of the individual channel cross sections were derived in ODDK22  the main innovation of this paper lies in the integration of these channels into a complete, data-driven model for the inclusive neutrino production cross section, along with a systematic quantification of the associated uncertainties. We work with neutrino energies from $10$ GeV to $10^{7}$ GeV, and correspondingly to incident proton (nuclei) energies from $10$ GeV to $10^{9}$ GeV (GeV/n). Our results include a realistic and conservative estimate of the uncertainties affecting the differential cross section $d\sigma/dE_{\nu}$, understood as the sum over all production channels. The uncertainty is estimated to be $5\%$ for $E_{\nu}\leq 1000$ GeV, increasing to $10\%$ at higher energies in the TeV range and reaching 20% at $E_{\nu}\simeq 10^{7}$ GeV.
@@ Acknowledgments
Appendix A Lorentz invariant cross sections
For the sake of completeness, we report all the formulas taken from ODDK22 and employed in this paper.
A.1 $\pi^{\pm}$ model
The Lorentz invariant cross section for $\pi^{\pm}$ is given by:
$\sigma_{\rm inv}=\sigma_{0}(s)\,c_{1}\,\Big[F_{p}(s,p_{T},x_{R})+F_{r}(p_{T},x_{R})\Big]\,A(s),$ (6)
where $\sigma_{0}(s)$ is the total inelastic $p+p$ cross section taken from ODDK22. The functional form of $F_{p}(p_{T},x_{R})$ is:
$F_{p}$ $(s,p_{T},x_{R})=(1-x_{R})^{c_{2}}\exp(-c_{3}\,x_{R})\,p_{T}^{c_{4}}$
$\times\exp\left[-c_{5}\sqrt{s/s_{0}}^{\;c_{6}}\left(\sqrt{p_{T}^{2}+m_{\pi}^{2}}-m_{\pi}\right)^{c_{7}\sqrt{s/s_{0}}^{\;c_{6}}}\right]\;,$
where $\sqrt{s_{0}}=17.3$ GeV. The model parameters $c_{i}$ are reported in Table II of ODDK22.
On the other hand, the functional form of $F_{r}(p_{T},x_{R})$ reads:
$F_{r}$ $(p_{T},x_{R})=(1-x_{R})^{c_{8}}$
$\times\exp\left[-c_{9}\,p_{T}-\left(\frac{|p_{T}-c_{10}|}{c_{11}}\right)^{c_{12}}\right]$
$\times\biggl[c_{13}\exp(-c_{14}\,p_{T}^{c_{15}}x_{R})+$
$\qquad+c_{16}\exp\left(-\left(\frac{|x_{R}-c_{17}|}{c_{18}}\right)^{c_{19}}\right)\biggr].$
Finally, we allow for an additional scaling with $\sqrt{s}$, which is required to obtain the correct $\pi^{\pm}$ multiplicity at different energies. The functional form is given by
$A(s)$ $=$ $\frac{1+\left(\sqrt{s/c_{20}}\right)^{c_{21}-c_{22}}}{1+\left(\sqrt{s_{0}/c_{20}}\right)^{\,c_{21}-c_{22}}}\,\left(\sqrt{\frac{s}{s_{0}}}\right)^{c_{22}}$ (9)
In all these formulas, $p_{T},\sqrt{s}$, the mass of the particles, $\sqrt{s_{0}}$ and energies are intrinsically normalized to 1 GeV, in order to have dimensionless parameters.
A.2 $K^{\pm}$ model
We define the Lorenz invariant cross section by:
$\sigma_{\rm inv}=\sigma_{0}(s)\,d_{1}\,F_{K}(s,p_{T},x_{R})\,A_{K}(s)$ (10)
with
$\!\!F_{K}(s,p_{T},x_{R})=(1-x_{R})^{d_{2}}\exp(-d_{3}\,p_{T}^{d_{4}}x_{R})\,p_{T}^{d_{5}}$ (11)
$\;\;\times\exp\left[-d_{6}\sqrt{s/s_{0}}^{\;d_{7}}\left(\sqrt{p_{T}^{2}+m_{K}^{2}}-m_{K}\right)^{d_{8}\sqrt{s/s_{0}}^{\;d_{7}}}\right]\;,$
where $m_{K}$ is the mass of the kaon and $d_{i}$ are the fit parameters reported in Table III of ODDK22. The energy dependent normalization $A_{K}(s)$ is taken to be:
$A_{K}(s)$ $=$ $A_{K}^{0}\!(1-\frac{\sqrt{s_{\rm th}}}{\sqrt{s}})\!(1+\sqrt{\frac{s}{d_{9}}}^{d_{10}-d_{11}})\!\sqrt{s}^{d_{11}}\;\;\;$ (12)
where $s_{\rm th}$ is the threshold energy for $K^{+}$ production and $A_{K}^{0}$ is determined by the condition $A_{K}(s_{0})=1$.
A.3 $K^{0}_{S}$ and $\Lambda$ model
We define the Lorenz invariant cross section of $K^{0}_{S}$ by:
$\sigma_{\rm inv}=\sigma_{0}(s)\,k_{1}\,F_{K^{0}_{S}}(p_{T},x_{F})\,A_{K^{0}_{S}}(s),$ (13)
with
$\!\!F_{K^{0}_{S}}(p_{T},x_{F})=(1-|x_{F}|)^{k_{2}}\times$ (14)
$\times\exp(-k_{3}\,p_{T}^{k_{4}}|x_{F}|)\;\;p_{T}^{k_{5}}\exp\left[-k_{6}\;p_{T}^{k_{7}}\right]\;,$
where $k_{i}$ are the fit parameters. The energy dependent normalization $A_{K^{0}_{S}}(s)$ is taken to be:
$A_{K^{0}_{S}}(s)$ $=$ $A_{K^{0}_{S},0}\left(1-\sqrt{\frac{k_{8}}{s}}^{k_{9}-k_{10}}\right)\sqrt{s}^{k_{10}},$ (15)
where the $A_{K^{0}_{S},0}$ is determined by the condition $A_{K^{0}_{S}}(\sqrt{s_{0}}=17.3\rm{\,GeV})=1$ and the best-fit parameters are reported in Table VII of ODDK22. The Lorentz-invariant cross section for $\Lambda$ follows the same functional form as that of $K^{0}_{S}$, differing only in the values of the fit parameters.